OpenStudy (anonymous):

identify the parent function from its function rule g(x)=(x-6)^2

OpenStudy (anonymous):

g(x) =y y = (x-6)^2 (y+0)=(x-6)^2 graphically, y can be translated as 0 and x as -6 so,the parent function is y=x^2